From GlueXWiki
Jump to: navigation, search

April 12, 2012 FDC meeting


  1. Production Construction Tracking (Dave)
    • Production status
  2. Results from the oxygen studies (Lubomir)
    • First production cell with Viton O-rings
    • New tests with EPDM
  3. Engineering update (Bill)
  4. Electronics update (Chris)
  5. Other


Participants: Bill, Dave, Chris, Nick, Mark, Simon, Beni, Eugene, and Lubomir.


- Dave: working on cathodes #44 and #45 for the fourth package, tension measurements today on wire plane #25 (second for fourth package), three wire planes waiting for deadening. The foil of the end window #10 (that was wavy) was removed: it turned out, by unknown reasons, the epoxy didn't cure after many days and the surface was still tacky. New foil will be tensioned.

Results from the oxygen studies

- Starting from configuration 12 in the table linked above, we are performing oxygen measurements at Blue Crab; Brooks gas system with two flow controllers and the oxygen sensor were moved there. Before looking into the results there was a long discussion about the way the measurements are done. We estimate the relative uncertainties of the measurements to be ~200 ppm, in addition we may have an overall factor of up to 30%.

- The first production cell with new Viton O-rings was installed and oxygen was measured (it took 4 days) 500-600 ppm at 330 ccpm gas flow and 900-1000 ppm at 220 ccpm. For comparison, the four cells with the original EPDM O-rings had ~7,000 ppm at 330 ccpm gas flow, so the new results show an improvement of more than a factor of 3, but a factor of ~2 worse than in the testing chamber. There was a significant leak, that later turned out to be caused by a small piece of Kapton tape on the top of the O-ring. Since yesterday, we replaced the top end window and cathode type-3 with a cathode type-1 and Lexan sheet. This is to investigate to role of the flatness and possible frame deformation.

- Due to the difficulties to maintain clean and flat surfaces needed for the Viton O-rings, we started looking again at the possibilities to use EPDM. We set up at Blue Crab a second oxygen testing place. There we started with the simplest configuration: spacer ring with EPDM O-rings between two Lexan plates: 220 ppm oxygen (to be compared to 400 ppm we measured before). As a next step, yesterday we added to the same configuration cathode type-1 with EPDM O-ring, expecting to have the final results tomorrow.

- Even we had a factor of 3 improvement, we expect ~0.5% oxygen for the whole package at 220ccpm. We would need a factor of 5 higher flow to have the oxygen below 0.1% what we had in the prototype. Another way to reduce the effect of the oxygen is to use less CO2, either going to 90/10 gas mixture or using methane instead of (some part of) CO2. Eugene: using organic gases may cause aging, but in Hall A where the rates are much higher they have used ethane. In any case, Eugene, Bill, Beni insisted that we don't want to be constrained by the gas and we should continue to fight for lower oxygen. Bill proposed to investigate the use of Apiezon (vacuum grease) on the O-rings. We will try it tomorrow with the EPDM O-rings on the spacer ring.


- Bill did Ansys analyses of the frame deformations with Viton O-rings (see plots attached above). The maximum frame deformation is less than 3mils, this is to be compared to 10 mils compression. It means in case of flat surfaces it will always seal the gas, however the compression will vary 30% or more in case of uneven surfaces.

- Bill is preparing a presentation for the fiducial scheme.


- Chris is ready to start installing the grounding on the first package. After the meeting we were at 126 looking into this. The parts for the ground cable connections have come and Beni will try them on the first package.